OpenStudy (anonymous):

Find Limit as X approaches to positive/negative infinity. y = (2-x/(x+1))(x^2/(5+x^2))

OpenStudy (anonymous):

\[y= \left( 2-\frac{ x }{ x+1 } \right)\left( \frac{ x^2 }{ 5+x^2 } \right)\]

ganeshie8 (ganeshie8):

divide x^2 in last fraction

ganeshie8 (ganeshie8):

\(\large y= \left( 2-\frac{ x }{ x+1 } \right)\left( \frac{ 1 }{ 5/x^2+1 } \right)\)

ganeshie8 (ganeshie8):

divide x in the first fraction as well

ganeshie8 (ganeshie8):

\(\large y= \left( 2-\frac{ 1 }{ 1+1/x } \right)\left( \frac{ 1 }{ 5/x^2+1 } \right)\)

ganeshie8 (ganeshie8):

now u may sub